Given a function described as equation y = 3x + 4, what is y when x is 1, 2, and 3?


7, 8, 9


3, 7, 11


3, 6, 9


7, 10, 13
1
Expert's answer
2020-02-27T10:54:02-0500

Solution:


y1 =y(1)= (3*1) + 4 = 7,

y2 =y(2)= (3*2) + 4 = 10,

y3 = y(3)=(3*3) + 4 = 13.


Answer: 7,10,13.